The angle that the string of a long pendulum makesThe angle that the string of a long pendulum makes with the vertical is shown as a function of time.

a) What is the angular frequency of the pendulum?

b)What is the amplitude of the pendulum's motion, in meters?
Answer #1

angular frequency = change in anglular position / time

3 degree = 0.0523599 radian

-3 degree = -0.0872664626 radian

angular frequency = 0.0523599 - (-0.0872664626) / 1

angular frequency = 0.0523599 - (-0.0872664626) / 1

angular frequency = 0.1396263626 rad/sec

2 * pi * sqrt(length / g) = 2 * pi / angular frequency

sqrt(length / 9.8) = 1 / 0.1396263626

length = 502.68 m

amplitude = length * sin(5)

amplitude = 502.68 * sin(5)

amplitude = 43.8114 m